(Size effect) Montebano, Inc. plans to choose between two manually exclusive projects. The first one, A, is a larger project costing $2,500,000 initially and generates $270,000 each year. Project A is expected to last 10 years. A smaller project, B has an initial cost of $300,000 and is expected to produce free cash flows of $50,000 per year over 10 years. Montebano has a weighted average cost of capital of 12%.

a. Calculate the NPV, and IRR.

b. Which project should be accepted?

c. Discuss the facts of this problem and the effects of Montebano's decision.
Answer #1
Year Project A PV @12% PV PV @1% PV Project B PV @12% PV PV @1% PV
1 270000 0.8929 241071 0.9901 267327 50000 0.8929 44643 0.9901 49505
2 270000 0.7972 215242 0.9803 264680 50000 0.7972 39860 0.9803 49015
3 270000 0.7118 192181 0.9706 262059 50000 0.7118 35589 0.9706 48530
4 270000 0.6355 171590 0.9610 259465 50000 0.6355 31776 0.9610 48049
5 270000 0.5674 153205 0.9515 256896 50000 0.5674 28371 0.9515 47573
6 270000 0.5066 136790 0.9420 254352 50000 0.5066 25332 0.9420 47102
7 270000 0.4523 122134 0.9327 251834 50000 0.4523 22617 0.9327 46636
8 270000 0.4039 109048 0.9235 249340 50000 0.4039 20194 0.9235 46174
9 270000 0.3606 97365 0.9143 246872 50000 0.3606 18031 0.9143 45717
10 270000 0.3220 86933 0.9053 244427 50000 0.3220 16099 0.9053 45264
Total Inflow 1525560 2557252 282511 473565
Intial Outflow 2500000 2500000 300000 300000
a) NPV (A-B) -974440 57252 -17489 173565
NPVa @ 1% 57252 NPVa @ 1% 173565
NPVb @ 12% -974440 NPVb @ 12% -17489
Ra 1.00% Ra 1.00%
Rb 12.00% Rb 12.00%
IRR = Ra + (NPVa * (Rb - Ra)) / NPVa- NPVb
IRR = 1% + ( 57252 * (12%-1%)) / 57252- (-974440) IRR = 1% + (173565 * (12%-1%)) / 173565- (-17489)
IRR = 1% + ( 57252 * 11% / 57252- (-974440) IRR = 1% + (173565 * 11% / 173565- (-17489)
IRR = 1% + ( 57252 * 11% / 57252+974440 IRR = 1% + (173565 * 11% / 191054
IRR = 1% + 6297 / 1031692 IRR = 1% + 19092 / 191054
IRR = 1% + 0.006 IRR = 1% + 0.099
a) IRR = 1.006% IRR = 1.099%
b) Non of the project is viable as NPV is negative in both the cases and IRR is less than cost of capital …i.e. 12%
c) Project initial cost is two high as comparative to PV inflow in 10 years for both of the cases
Projected cost of capital is high as compare to IRR for both the cases
The Scarp Value from both the project is NIL
The project is not acceptable as both the project creating wealth